OpenStudy (anonymous):

Solve the system of equations below by graphing both equations with a pencil and paper. What is the solution? y = x - 5 y = -2x + 4

OpenStudy (anonymous):

how would the bottom look?

OpenStudy (anonymous):

Ok so we have... x-y=5 on top and 2x+y=4 on the bottom Multiply the top by -2 and we have: -2x+2y=-10 2x+y=4 on the bottom the x's cancel out leaving you with 3y=-6 divide by 3 and you have y=-2 plug back in and you will find that x = 3